# IMAIA: Interactive Maps AI Assistant for Travel Planning and Geo-Spatial Intelligence

### v3

@@ IV-D Evaluation on Spatial Understanding
V Conclusion
In this work, we introduced IMAIA, an interactive Maps AI Assistant that unifies language, maps, and geospatial signals to enable natural, spatially grounded interaction. Through its two complementary components—Maps Plus for map-centric question answering and PAISA for camera-aware place understanding—IMAIA significantly improves the accuracy, responsiveness, and usability of geospatial reasoning compared to strong baselines. Our experiments show that Maps Plus boosts place detection accuracy from under 43% with conventional approaches to nearly 90% by treating maps as first-class context and incorporating structured grounding signals. For navigation, PAISA demonstrates that human-centered, direction-based guidance aligns more closely with natural pedestrian behavior than rigid turn-by-turn instructions, offering both flexibility and reduced detours. Beyond place recognition and navigation, our distilled spatial reasoning model achieves 84% accuracy and a 7.3 inference speedup over agent-based pipelines, highlighting the effectiveness of lightweight task-aligned distillation for real-time deployment. Together, these results demonstrate that IMAIA represents a step toward conversational mapping that is both accurate and practical. By integrating visual, spatial, and linguistic cues into a unified framework, it enables richer scene understanding, more intuitive navigation, and scalable performance for user-facing applications.

### v4

@@ 4.4 Evaluation on Spatial Understanding
5 Conclusion
In this work, we introduce IMAIA, an interactive Maps AI Assistant that spans desktop-style map exploration and on-the-ground, last–100–meter navigation in a unified, deployment-oriented framework. IMAIA is built around two complementary components—Maps Plus for map-centric spatial understanding and PAISA for ego-centric scene understanding and geospatial place grounding—coordinated by a lightweight multi-agent orchestration layer. Maps Plus treats the map viewport as a quadkey-indexed visual prompt, enabling view-conditioned, map-centric queries and boosting place detection accuracy from under 43% with conventional approaches to nearly 90%. PAISA operates in an AR-style setting, fusing camera input with geospatial signals to provide human-centered, bearing-aware guidance that better matches natural pedestrian behavior than rigid turn-by-turn instructions, reducing unnecessary detours. Beyond individual modules, our results show that IMAIA’s model-agnostic, modular design supports efficient and practical geospatial reasoning for user-facing applications. A distilled spatial reasoning model within the framework attains 84% accuracy and a 7.3 inference speedup over agent-based pipelines (1.7s vs. 12.4s), demonstrating the effectiveness of lightweight, task-aligned distillation. Together, these findings indicate that focusing on system design–quadkey-based visual prompting for maps, AR-style camera grounding, and a replaceable VLM/vision backend enables an interactive and responsive geospatial experience and production-ready quality for trip planning and local discovery.